# Fine-Tuning Universal Machine-Learned Interatomic Potentials: A Tutorial on Methods and Applications

### v1

@@ Abstract
Universal machine-learned interatomic potentials (U-MLIPs) have demonstrated broad applicability across diverse atomistic systems but often require fine-tuning to achieve task-specific accuracy. While the number of available U-MLIPs and their fine-tuning applications is rapidly expanding, there remains a lack of systematic guidance on how to effectively fine-tune these models. This tutorial provides a comprehensive, step-by-step guide to fine-tuning U-MLIPs for computational materials modeling. Using the recently released MACE-MP-0 as a representative example, we cover key aspects of the fine-tuning process, including dataset preparation, hyperparameter selection, model training, and validation. The effectiveness of fine-tuning is demonstrated through case studies involving force prediction in solid-state electrolytes, stacking fault defects in metals, and solid–liquid interfacial interactions in low-dimensional systems. To support practical applications, we include code examples that enable researchers, particularly those new to the field, to efficiently incorporate fine-tuned U-MLIPs into their workflows.
@@ I INTRODUCTION
Table 1: List of the representative U-MLIPs. See Appendix V.1 for further details.
@@
While MACE-MP-0 and its varaints have demonstrated robust performance in molecular dynamics simulations [22], its predictive accuracy—like that of other U-MLIPs—remains limited for certain tasks, particularly in modeling mechanical properties such as elastic constants and stacking fault energies in elemental alloys [39]. This underscores the need for fine-tuning, which adapts pre-trained models to specific systems by refining their parameters using high-fidelity computational or experimental data. Although the fine-tuning of atomistic foundation models has garnered growing attention, comprehensive and systematic studies remain limited, especially practical tutorials that guide beginners through the fine-tuning process. Recent works have started addressing this gap [40, 41, 42, 43, 44, 45, 46, 47, 48], and continued efforts are expected to clarify best practices across various application scenarios.
@@
Beyond the pursuit of higher accuracy, a central motivation for fine-tuning in the context of machine-learned interatomic potentials lies in a fundamental question: given a fixed dataset, should one fine-tune an existing foundation model or train a new model from scratch? In practice, fine-tuning is often more efficient in terms of training time and data usage, and in many cases yields comparable or even superior accuracy. While foundation models, pre-trained on large and diverse datasets, effectively capture general chemical and structural patterns, they often lack the precision required for specific systems. Fine-tuning addresses this limitation by incorporating system-specific data, thereby enhancing accuracy with significantly lower data and computational costs. Moreover, fine-tuned models typically converge faster due to the advantage of informed initialization. Therefore, when a suitable foundation model is available, fine-tuning presents a compelling alternative to training-from-scratch.
@@
This tutorial offers a practical guide to fine-tuning U-MLIPs, with a particular focus on helping researchers who are new to the area. The goal is to provide step-by-step instructions for preparing data, training models, selecting hyperparameters, and applying the resulting potentials in various atomistic simulations. To keep the presentation accessible, theoretical details are kept to a minimum, with references provided for readers seeking a deeper mathematical and physical understanding. The methods introduced here are demonstrated using widely used tools such as ASE, LAMMPS, and the Random Batch Molecular Dynamics (RBMD) interface [49]. MACE-MP-0 is selected as the main example due to its broad applicability, high accuracy across benchmark tasks, and demonstrated stability in molecular dynamics simulations [22]. While this tutorial is based on MACE-MP-0, the fine-tuning workflow and practical considerations apply more generally to other U-MLIPs as well  [^1] A note of caution: The procedures and recommendations in this tutorial reflect the state of the field at the time of writing. Given the rapid pace of developments in model architectures, datasets, and fine-tuning strategies, updates are likely in the near future..
@@
It is worth noting that the fine-tuning of U-MLIPs can be efficiently carried out using the RBMD package, which enables large-scale particle simulations at the nano- and microscale. In contrast to conventional molecular dynamics frameworks, RBMD leverages random batch algorithms [51, 52, 53, 54] to efficiently handle nonbonded interactions, supporting simulations of up to 10 million particles on a single GPU with a CPU core [49]. Moving forward, we aim to integrate the RBMD platform with MACE-based MLIPs for more practical and large-scale simulations. To support this integration, an open-access platform is available at https://www.randbatch.com/rbmd, where further examples and implementation scripts for realistic simulation scenarios will be released in the future.
@@
The remainder of this tutorial is organized as follows. In Section II, we introduce the overall fine-tuning workflow for U-MLIPs, including a brief overview of the MACE-MP-0 framework, data preparation, potential training, and model validation. Section III presents three illustrative applications of fine-tuning U-MLIPs: (i) force accuracy evaluation in the solid-state electrolyte Li10GeP2S12 (LGPS), (ii) stacking fault energetics in body-centered cubic molybdenum (Mo), and (iii) surface interactions at the graphene–water interface. Finally, Section IV provides a summary and outlook on the development and fine-tuning of U-MLIP models.
@@ II METHODS
The fine-tuning procedure is supported by the official MACE repository: https://github.com/ACEsuit/mace. Installation instructions are also available in the tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git. For completeness, Table 8 in Appendix V.3 provides an overview of relevant repositories and their functionalities. For instance, the RBMD interface [49] supports both fine-tuning and molecular dynamics simulations with MACE-based MLIPs.
@@ II.1 The MACE-MP-0 Model: Architecture, Fine-Tuning, and Recent Developments
We begin by reviewing the MACE-MP-0 model in Section II.1.1. The fine-tuning procedures introduced here are broadly applicable to other U-MLIPs with only minor adjustments. For completeness, a brief overview of other representative U-MLIP models is provided in Appendix V.1.
@@ II.1.1 Overview of the MACE-MP-0 Model
Before introducing U-MLIPs, we briefly recall the concept of MLIPs. These models map atomic positions and chemical elements to the potential energy of a given atomic system. The MACE architecture [6] is a representative example, extending the Atomic Cluster Expansion (ACE) framework [55, 5, 37], which was originally developed for MLIPs and has since found applications beyond atomistic modeling [56, 57, 58, 59]. MACE is an equivariant message-passing graph tensor network [60, 61], where many-body atomic geometry is encoded layer by layer. Messages are built from linear combinations of tensor product bases derived from two-body permutation-invariant polynomials in a spherical basis [5, 37]. Equivariance is maintained through tensor contractions with irreducible representations and generalized Clebsch–Gordan coefficients [62, 38]. The final output represents the per-atom contribution to the total potential energy. Further architectural details are provided in Appendix V.2 and in Ref. [6].
@@
Multi-head fine-tuning extends this idea by attaching multiple task-specific output heads to a shared pre-trained backbone [69]. Each head is tailored to a different target system or property, while the backbone captures general representations. By partially freezing the backbone, this approach enables efficient training across multiple tasks and reduces redundancy. An implementation of this method is available at https://github.com/ACEsuit/mace/tree/main.
II.1.2 Recent Updates
@@ II.2 Data Preparation
A key question in the fine-tuning of U-MLIPs is how to select a dataset that leads to better performance—both in terms of accuracy and data efficiency—compared to using the original foundation model or training a new model from scratch. In this section, we summarize a range of strategies, from empirical approaches based on physical intuition (Section II.2.1) to more automated data selection guided by uncertainty quantification (Section II.2.2).
@@ II.2.2 Uncertainty-Based Dataset Construction
In cases where a large pool of configurations is available (e.g., from long MD simulations or existing datasets), uncertainty-based filtering can be applied to construct an efficient fine-tuning set. Here, an uncertainty estimator (such as an ensemble or a surrogate model with Bayesian regression) is used to assign a confidence value to each configuration. Configurations with higher estimated uncertainty are prioritized, as they are more likely to contribute new information to the model. This strategy has been used successfully in recent works, such as Ref. [77], where a linear ACE model is trained on top of an existing representation to guide uncertainty filtering.
@@ II.3 Fine-Tuning
Based on our experience, we suggest using at least $L=1$ (i.e. medium) models for a balance between accuracy and efficiency. The following recommendations are in general helpful on stability and accuracy of the resulting model from fine-tuning:
@@
  • Number of samples in the pretrained head (num_samples_pt): A larger value helps preserve more information and improves stability in molecular dynamics. However, there is a trade-off: an overly large value may reduce the achievable accuracy.
We recommend monitoring convergence behavior (cf. Fig. 2) and validation accuracy throughout the fine-tuning process. Evaluating model performance on unseen configurations is crucial and will be discussed in the next subsection.
@@ II.4 Model Validation and Evaluation
We refer readers to our third case study (Section III.3) for a complete example of running MD simulations in LAMMPS using a fine-tuned model. All input files and scripts are available in our public tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
@@ III EXAMPLES AND APPLICATIONS
The fine-tuned models integrate seamlessly with widely used atomistic simulation platforms such as ASE, LAMMPS and RBMD (see Appendix V.3), enabling efficient molecular dynamics and other atomistic simulations. In this section, we illustrate the effectiveness of fine-tuning using representative case studies on solid-state electrolytes, metallic stacking faults, and interfacial phenomena in low-dimensional systems. The pretrained foundation models are publicly available at https://github.com/ACEsuit/mace-mp/. In addition, we provide fine-tuning scripts, example outputs, and best practices in our tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
@@ III.1 Li10GeP2S12
In this example, we demonstrate the fine-tuning procedure on a solid-state electrolyte system, lithium germanium phosphorous sulfide (Li10GeP2S12, abbreviated as LGPS), using a dataset published on DPAISquare [73]. LGPS is a prototypical member of the argyrodite family and has attracted significant attention due to its high lithium-ion conductivity, making it a promising candidate for next-generation all-solid-state batteries. Modeling such materials requires accurate interatomic potentials capable of describing both local structural features and ionic transport mechanisms, making LGPS an ideal testbed for evaluating the transferability and fine-tuning of U-MLIPs.
@@
Table 4: Test accuracy of different fine-tuned MACE foundation models, reported as RMSE of energy (E) and force (F), relative to DFT values. “Relative F” denotes the force RMSE normalized by the standard deviation of DFT forces. Each model is evaluated before and after SWA, with the epoch number indicating the training step.
@@
MACE-0b3 (110) 0.32 16.15 1.92
MACE-0b3 (148) 0.28 15.84 1.87
MACE-mpa (110) 0.30 15.07 1.79
MACE-mpa (146) 0.27 14.88 1.77
MACE-omat (110) 0.35 15.46 1.84
MACE-omat (149) 0.27 15.32 1.82
Figure 4 compares the test accuracy of models fine-tuned from different MACE foundation models with that of training-from-scratch, evaluated at different fractions of the training dataset (10%, 50%, and 100%). Figure 4(a) shows the RMSE in energy, while Figure 4(b) shows the RMSE in forces. Across all dataset sizes, fine-tuned models (mp-0b3, mpa-0, omat-0) consistently achieve lower errors than training-from-scratch, particularly at lower data fractions, demonstrating the data efficiency and accuracy benefits of fine-tuning. The results highlight the robustness of MACE foundation models as effective starting points for fine-tuning, leading to improved performance with reduced training data.
Figure 4: Comparison between fine-tuning and training-from-scratch in terms of data efficiency: (a) RMSE in energy (meV/atom) and (b) RMSE in forces (meV/Å).
III.2 Molybdenum

### v2

@@ Abstract
Universal machine-learned interatomic potentials (U-MLIPs) have demonstrated broad applicability across diverse atomistic systems but often require fine-tuning to achieve task-specific accuracy. While the number of available U-MLIPs and their fine-tuning applications is rapidly expanding, there remains a lack of systematic guidance on how to effectively fine-tune these models. This tutorial provides a comprehensive, step-by-step guide to fine-tuning U-MLIPs for computational materials modeling. Using the recently released MACE-MP-0 as a representative foundation model, we illustrate the full workflow of dataset preparation, hyperparameter selection, model training, and validation. Beyond methodological guidance, we conduct systematic case studies on solid-state electrolytes, stacking fault defects in metals, semiconductors, solid–liquid interfacial interactions in low-dimensional systems, and more complicated heterointerfaces. These examples demonstrate that fine-tuning substantially improves predictive accuracy while maintaining affordable computational cost, accelerates training convergence, enhances out-of-distribution generalization, and achieves superior data efficiency. Remarkably, fine-tuned foundation models can even capture aspects of long-range physics without explicit corrections. Together, these results highlight that fine-tuning not only provides a practical recipe for applying U-MLIPs, but also offers new insights into their physical fidelity and potential for advancing large-scale atomistic simulations. To support practical applications, we include code examples that enable researchers, particularly those new to the field, to efficiently incorporate fine-tuned U-MLIPs into their workflows.
@@ I INTRODUCTION
Table 1: List of the representative U-MLIPs.
@@
While MACE-MP-0 and its varaints have demonstrated robust performance in molecular dynamics simulations [22], its predictive accuracy—like that of other U-MLIPs—remains limited for certain tasks, particularly in modeling mechanical properties such as elastic constants and stacking fault energies in elemental alloys [39]. This underscores the need for fine-tuning, which adapts pre-trained models to specific systems by refining their parameters using high-fidelity computational or experimental data. Although the fine-tuning of atomistic foundation models has garnered growing attention, comprehensive and systematic studies remain limited, especially practical tutorials that guide beginners through the fine-tuning process. Recent works have started addressing this gap [40, 41, 42, 43, 44, 45, 46, 47, 48], and continued efforts are expected to clarify best practices across various application scenarios. While the number of available U-MLIPs and their fine-tuning applications is rapidly expanding, there remains a lack of systematic guidance on how to effectively fine-tune these models, which serves as the primary motivation of this tutorial.
@@
Beyond the pursuit of higher accuracy, a central motivation for fine-tuning in the context of MLIPs lies in a fundamental question: given a fixed dataset, should one fine-tune an existing foundation model or train a new model from scratch? In practice, fine-tuning is often more efficient in terms of training time and data usage, and in many cases yields comparable or even superior accuracy. While foundation models, pre-trained on large and diverse datasets, effectively capture general chemical and structural patterns, they often lack the precision required for specific systems. Fine-tuning addresses this limitation by incorporating system-specific data, thereby enhancing accuracy with significantly lower data and computational costs. Moreover, fine-tuned models typically converge faster due to the advantage of informed initialization. Therefore, when a suitable foundation model is available, fine-tuning presents a compelling alternative to training-from-scratch.
@@
This tutorial provides a practical guide to fine-tuning U-MLIPs, aiming to support researchers—especially those new to the field—by filling the current gap in systematic guidance for fine-tuning foundation models. The goal is to provide step-by-step instructions for preparing data, training models, selecting hyperparameters, and applying the resulting potentials in various atomistic simulations. To keep the presentation accessible, theoretical details are kept to a minimum, with references provided for readers seeking a deeper mathematical and physical understanding. The methods introduced here are demonstrated using widely used tools such as ASE, LAMMPS, and the Random Batch Molecular Dynamics (RBMD) interface [49]. MACE-MP-0 is selected as the main example due to its broad applicability, high accuracy across benchmark tasks, and demonstrated stability in molecular dynamics simulations [22].
Beyond methodological guidance, we conduct extensive numerical experiments that yield several key insights. Fine-tuning consistently improves predictive accuracy and accelerates convergence while maintaining affordable computational cost. It enhances out-of-distribution generalization and achieves higher data efficiency compared to training from scratch, even when using reduced datasets. Interestingly, fine-tuned foundation models sometimes show improved accuracy even for cases involving long-range effects, suggesting that certain aspects of such interactions may already be partially captured during pre-training. Taken together, these findings show that fine-tuning not only offers a practical recipe for applying U-MLIPs, but also provides deeper insights into their physical fidelity and their potential to advance large-scale atomistic simulations.
@@
It is worth noting that the fine-tuning of MACE-MP-0 can be efficiently carried out using the RBMD package, which enables large-scale particle simulations at the nano- and microscale. In contrast to conventional molecular dynamics frameworks, RBMD leverages random batch algorithms [50, 51, 52, 53] to efficiently handle nonbonded interactions, supporting simulations of up to 10 million particles on a single GPU with a CPU core [49]. Moving forward, we aim to integrate the RBMD platform with MACE-based MLIPs for more practical and large-scale simulations. To support this integration, an open-access platform is available at https://www.randbatch.com/rbmd, where further examples and implementation scripts for realistic simulation scenarios will be released in the future.
@@
The remainder of this tutorial is organized as follows. In Section II, we introduce the overall fine-tuning workflow for U-MLIPs, including a brief overview of the MACE-MP-0 framework, data preparation, potential training, and model validation. Section III presents six illustrative applications of fine-tuning U-MLIPs: (i) force accuracy evaluation in the solid-state electrolyte Li10GeP2S12 (LGPS), (ii) stacking fault energetics in body-centered cubic molybdenum (Mo), (iii) generalization tests on out-of-distribution in semicondutor silicon, (iv) surface interactions at the graphene–water interface, (v) molecule–oxide interactions, and (vi) solid–solid interfaces, the latter two designed to assess the ability of fine-tuned models to capture long-range effects. Finally, Section IV provides a summary and outlook on the development and fine-tuning of U-MLIP models.
@@ II METHODS
The fine-tuning procedure is supported by the official MACE repository: https://github.com/ACEsuit/mace. Installation instructions are also available in the tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git. The RBMD interface [49] supports both fine-tuning and molecular dynamics simulations with MACE-based MLIPs.
@@ II.1 The MACE-MP-0 Model: Architecture, Fine-Tuning, and Recent Developments
We begin by reviewing the MACE-MP-0 model in Section II.1.1. The fine-tuning procedures introduced here are broadly applicable to other U-MLIPs with only minor adjustments.
@@ II.1.1 Overview of the MACE-MP-0 Model
Before introducing U-MLIPs, we briefly recall the concept of MLIPs. These models map atomic positions and chemical elements to the potential energy of a given atomic system. The MACE architecture [6] is a representative example, extending the Atomic Cluster Expansion (ACE) framework [54, 5, 37], which was originally developed for MLIPs and has since found applications beyond atomistic modeling [55, 56, 57, 58]. MACE is an equivariant message-passing graph tensor network [59, 60], where many-body atomic geometry is encoded layer by layer. Messages are built from linear combinations of tensor product bases derived from two-body permutation-invariant polynomials in a spherical basis [5, 37]. Equivariance is maintained through tensor contractions with irreducible representations and generalized Clebsch–Gordan coefficients [61, 38]. The final output represents the per-atom contribution to the total potential energy. Further architectural details are provided in Ref. [6].
@@ II.1.2 Fine-Tuning Strategies
Multi-head replay fine-tuning extends this idea by attaching multiple task-specific output heads to a shared pre-trained backbone [68]. Each head is tailored to a different target system or property, while the backbone captures general representations. By partially freezing the backbone, this approach enables efficient training across multiple tasks and reduces redundancy. This design also mitigates catastrophic forgetting by preserving the performance of previously trained heads while adapting to new tasks. An implementation of this method is available at https://github.com/ACEsuit/mace/tree/main.
From an optimization perspective, existing fine-tuning strategies for U-MLIPs can be broadly categorized into: (i) Full-parameter fine-tuning, where all backbone and head parameters are updated [69]. This maximizes task-specific adaptability but incurs higher computational cost and greater risk of overfitting, especially when the target dataset is small. (ii) Partial freezing, where early layers or invariant feature extractors are kept fixed while only later layers or heads are updated. This reduces the number of trainable parameters, leading to faster convergence and better knowledge retention from the pre-trained model, but may limit adaptation to systems with strong distribution shifts [70]. (iii) Lightweight adaptation layers (e.g., low-rank adapters or bias tuning), which introduce a small set of trainable parameters on top of the frozen backbone [71]. This approach offers a favorable trade-off between efficiency and accuracy, and has shown promise in retaining generalization while improving data efficiency.
In molecular modeling, the choice of strategy should be guided by the similarity between the pre-training and target domains, the size and diversity of the fine-tuning dataset, and the computational budget. A comprehensive benchmark to determine the optimal strategy under different scenarios is still lacking, but future work will aim to provide such guidance.
II.1.3 Recent Updates
@@ II.2.2 Uncertainty-Based Dataset Construction
In cases where a large pool of configurations is available (e.g., from long MD simulations or existing datasets), uncertainty-based filtering can be applied to construct an efficient fine-tuning set. Here, an uncertainty estimator (such as an ensemble or a surrogate model with Bayesian regression) is used to assign a confidence value to each configuration. Configurations with higher estimated uncertainty are prioritized, as they are more likely to contribute new information to the model. This strategy has been used successfully in recent works, such as Ref. [79], where a linear ACE model is trained on top of an existing representation to guide uncertainty filtering. We evaluate the data efficiency of fine-tuning by comparing it against uncertainty-based filtering in Section III.2.
@@
For novice users, we recommend the following practical guidelines: (i) Small, chemically narrow systems: Start with a simple ensemble-based UQ (e.g., multiple MACE models trained from different random seeds) and select new configurations with the highest predicted force uncertainty. (ii) Chemically diverse or poorly explored systems: Consider dropout-based UQ during inference to reduce training cost while still capturing model variance, but apply a post-hoc calibration step such as isotonic regression or temperature scaling. (iii) When DFT labeling is expensive: Limit the number of configurations per iteration by combining UQ-driven selection with structural diversity filters (e.g., farthest-point sampling in descriptor space [107]). These approaches balance exploration efficiency, calibration accuracy, and computational feasibility, enabling active learning workflows that are both effective and accessible to new users.
@@ II.3 Fine-Tuning
From our experience, a medium-sized model ($L=1$) provides the most balanced trade-off between accuracy and computational efficiency. While the small model ($L=0$) is computationally inexpensive but lacks expressivity, the large model ($L=2$) increases accuracy at the cost of significantly reduced efficiency. The following recommendations are in general helpful on stability and accuracy of the resulting model from fine-tuning:
@@
  • Number of samples in the pretrained head (num_samples_pt): This parameter controls how many samples from the pretrained model’s output distribution are retained for the fine-tuning stage. A larger value preserves a richer set of features and variance from the pretrained head, which can improve MD stability. In other words, it acts as a form of “knowledge retention” that mitigates catastrophic forgetting of the pretrained model. However, there is a trade-off: if it is too large, the fine-tuned model remains overly biased towards the pretrained distribution, which may limit its ability to fit task-specific data and thus cap the achievable accuracy.
We recommend monitoring convergence behavior (cf. Fig. 2) and validation accuracy throughout the fine-tuning process. Evaluating model performance on unseen configurations is crucial and will be discussed in the next subsection.V
@@ II.4 Model Validation and Evaluation
We refer readers to our third case study (Section III.4) for a complete example of running MD simulations in LAMMPS using a fine-tuned model. All input files and scripts are available in our public tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
@@ III EXAMPLES AND APPLICATIONS
The fine-tuned models integrate seamlessly with widely used atomistic simulation platforms such as ASE, LAMMPS and RBMD (see Appendix VI), enabling efficient molecular dynamics and other atomistic simulations. In this section, we illustrate the effectiveness of fine-tuning using representative case studies on solid-state electrolytes, metallic stacking faults, and interfacial phenomena in low-dimensional systems. The pretrained foundation models are publicly available at https://github.com/ACEsuit/mace-mp/. In addition, we provide fine-tuning scripts, example outputs, and best practices in our tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
We performed fine-tuning and molecular dynamics simulations on a single NVIDIA A800 GPU with 80 GB of HBM memory. The software environment was configured with CUDA 11.7 and the NVIDIA HPC SDK (nvhpc) 20.9. Unless otherwise specified, all training runs were carried out under this single-GPU setup.
We present four representative examples to illustrate the fine-tuning of U-MLIPs across different materials contexts. Solid-state electrolytes are chosen to assess improvements in force prediction essential for accurate ion-transport simulations. Stacking fault defects in metals are included to evaluate generalization to defect configurations and the accurate prediction of stacking fault energies. Silicon with a diverse and complex dataset is used to evaluate out-of-distribution generalization. Solid–liquid interfacial systems are examined to test the model’s ability to capture interfacial energetics critical to phase transitions and growth processes.
@@ III.1 Li10GeP2S12
In this example, we demonstrate the fine-tuning procedure on a solid-state electrolyte system, lithium germanium phosphorous sulfide (Li10GeP2S12, abbreviated as LGPS), using a dataset published on DPAISquare [75]. LGPS is a prototypical member of the LGPS-type thiophosphate family, and has attracted significant attention due to its high lithium-ion conductivity, making it a promising candidate for next-generation all-solid-state batteries. Modeling such materials requires accurate interatomic potentials capable of describing both local structural features and ionic transport mechanisms, making LGPS an ideal testbed for evaluating the transferability and fine-tuning of U-MLIPs.
@@
Table 4: Test accuracy of different fine-tuned MACE foundation models, reported as RMSE of energy (E) and force (F), relative to DFT values. “Relative F” denotes the force RMSE normalized by the standard deviation of DFT forces. Each model is evaluated before and after SWA, with the epoch number indicating the training step. Fine-tuning consistently achieves good accuracy regardless of which foundation model is used.
@@
MACE-MP-0b3 (110) 0.32 16.15 1.92
MACE-MP-0b3 (148) 0.28 15.84 1.87
MACE-MPA-0 (110) 0.30 15.07 1.79
MACE-MPA-0 (146) 0.27 14.88 1.77
MACE-OMAT-0 (110) 0.35 15.46 1.84
MACE-OMAT-0 (149) 0.27 15.32 1.82
Figure 4 compares the test accuracy of models fine-tuned from different MACE foundation models with that of training-from-scratch, evaluated at different fractions of the training dataset (10%, 30%, 50%, 75% and 100%). Figure 4(a) shows the RMSE in energy, while Figure 4(b) shows the RMSE in forces. Across all dataset sizes, fine-tuned models (MP-0b3, MPA-0, OMAT-0) consistently achieve lower errors than training-from-scratch, particularly at lower data fractions, demonstrating the data efficiency and accuracy benefits of fine-tuning. The error bars remain small across experiments, indicating robustness against training noise. Overall, these results highlight the strength of MACE foundation models as effective starting points for fine-tuning, enabling improved performance even with limited training data.
Figure 4: Comparison between fine-tuning and training-from-scratch in terms of data efficiency: (a) RMSE in energy (meV/atom) and (b) RMSE in forces (meV/Å). Error bars represent one standard deviation, computed from three independent runs, with markers showing the corresponding mean values.
III.2 Silicon
We use silicon studied in [108] as a representative system to evaluate the out-of-distribution generalization of fine-tuning, since the potential risk of overfitting has attracted considerable attention in the fine-tuning of foundation models. The training set consists of standard bulk and defect configurations, while the test set includes grain boundaries, di-interstitials, stacking fault paths, and an amorphous configuration. Unlike a random split, the test set contains configurations entirely distinct from the training data, providing a stringent benchmark for assessing extrapolative performance.
The previous example highlighted the data efficiency of fine-tuning with limited training data. In this case, we focus on ensemble-uncertainty-based active learning, where candidate configurations are filtered according to prediction uncertainty (cf. Section II.2.2) following the procedure in [79]. This reduces the training set to only 5% of the original data. We then train a MACE model on this reduced dataset and compare its accuracy with that obtained from direct fine-tuning.
Figure 5 compares the performance of the foundation MACE model (MACE-MPA-0) with its fine-tuned counterpart (MACE-FT): (a) energy RMSE per atom (meV/atom), and (b) force RMSE (meV/Å). In both metrics, MACE-FT achieves substantially lower errors than MACE-MPA-0, indicating improved agreement with the reference data even for the out-of-distribution test set. This demonstrates that fine-tuning does not suffer from generalization issues in foundation models. A second conclusion is drawn from the comparison with models trained on uncertainty-filtered datasets (labeled “MACE-AL” in Figure 5). Although the filtered set contains a similar number of samples, fine-tuning still yields superior accuracy. This highlights that fine-tuning not only ensures generalization but also achieves better data efficiency than training from scratch on a reduced dataset obtained through UQ-based active learning.
Figure 5: Comparison of a foundation MACE model (MACE-MPA-0) and its fine-tuned counterpart (MACE-FT) on the silicon out-of-distribution test set. (a) Atomic structure; (b) energy RMSE per atom (meV/atom); (c) force RMSE (meV/Å). In both cases, MACE-FT achieves significantly lower errors than MACE-MPA-0, demonstrating improved accuracy and generalization. Moreover, MACE-FT also outperforms models trained from uncertainty-filtered datasets (MACE-AL).
III.3 Molybdenum